A wire 77 inches is cut into two pieces whose lengths have a ratio of 5 to 2. What is the length of each piece?
Answer by Alan3354(69443)

A wire 77 inches is cut into two pieces whose lengths have a ratio of 5 to 2. What is the length of each piece?
--------------
5 + 2 = 7
77/7 = 11
-----
5*11 = 55
2*11 = 22